A creditor provides a company with $100,000 cash by purchasing bonds at the end of Year 1. The creditor receives $6,000 in interest from the company during Year 2. The creditor sells the bonds at the end of Year 2 or $102,000. What was the rate of return generated by this investment? (Round the percentage to the nearest whole number.)
Answer #1
Rate of return = (C+(P1-P0))/P0 Where,
= (6000+(102000-100000))/100000 C = $             6,000
= 8% P0 = $       1,00,000
P1 = $       1,02,000